# Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks

### v2

@@ IV Conclusion
In this study, we developed and analyzed two hypergraph-based clustering algorithms, DIPHW and CoMem-DIPHW for improved cell clustering performance on scRNA-seq data. Firstly, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW) we developed extends the edge-dependent vertex weight (EDVW) hypergraph random walk proposed by [41], and consistently outperformed it in our experiments. DIPHW’s key idea lies in its symmetrized account of importance: it considers both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering.
Secondly, the CoMem and CoMem-DIPHW clustering methods we developed incorporate a Memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks (unipartite network projections of the scRNA-seq abundance counts: gene co-expression networks and cell co-expression networks). This approach allows us to integrate both local and global information when computing embeddings for clustering.
Thirdly, we identified and showed with experiments 3 the issue of inflated signals in co-expression networks brought by sparsity of scRNA-seq data, which thus provided additional motivation for conceptualizing the scRNA-seq data as hypergraphs for embedding computation. The sparsity-induced correlation inflation leads to misleading cell connections in cell co-expression networks. Conventional correlation measures (Pearson’s, Spearman’s, Kendall’s Tau) either treat the common zeros in cell expression profiles as a sign of homogeneity or ignore them during correlation computation (as in cosine similarity). Neither strategy handles the intricacy of the data sparsity well enough to avoid suboptimal results. We showed in Tab. 3 that unipartite network-based clustering approaches perform poorly on correlation networks constructed using cosine similarity compared to using other correlation measures. Furthermore, unipartite network-based clustering methods generally underperform compared to hypergraph-based methods or PCA, which do not require transforming scRNA-seq count data into unipartite networks.
Fourthly, we compared our proposed methods with well-established clustering methods commonly used in scRNA-seq analysis packages. These include five embedding-based methods (PCA, Node2Vec, EDVW [41], UMAP, and t-SNE) and four community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden). Our clustering performance comparison experiments span comprehensive parameter regimes, with CoMem-DIPHW and DIPHW consistently delivering superior performance, especially when data modularity is weak. In these weakly modular regimes, where clusters are hardly detectable, all the established methods we used for comparison show a much more pronounced decay in performance compared to CoMem-DIPHW. For empirical analysis, we tested our algorithms on 4 real scRNA-seq datasets and 2 benchmark datasets. In these experiments, CoMem-DIPHW consistently demonstrated superior performance in differentiating and classifying distinct cell types compared to the baseline methods.
Additionally, we designed a versatile simulation algorithm for scRNA-seq data that is highly customizable through its many user-specifiable parameters. This algorithm is designed to model a wide range of expression scenarios and modular structures. Notably, it incorporates intermodular co-expressions, which mimic cell-type cross-talk observed in real biological systems.
For future work, we plan to optimize the implementation of CoMem-DIPHW to reduce its memory complexity, enhancing its scalability for larger datasets. We are also interested in exploring alternative ways to incorporate the memory mechanism into the hypergraph random walk. Currently, CoMem employs a simple product approach, moving beyond the simple product used in CoMem could potentially uncover even more effective clustering strategies.
@@
.1 Embedding Visualizations
@@
Figure S7: CoMem-DIPHW Clustering Performance Assessment on the Human Brain Dataset [57] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S8: PCA Clustering Performance Assessment on the Human Brain Dataset [57] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
@@
Figure S9: CoMem-DIPHW Clustering Performance Assessment on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S10: PCA Clustering Performance Assessment on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S11: CoMem-DIPHW Clustering Performance Assessment on the Mouse Brain Dataset [59] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. (c) Cell type annotation using the PanglaoDB marker database.
Figure S12: PCA Clustering Performance Assessment on the Mouse Brain Dataset [59] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. (c) Cell type annotation using the PanglaoDB marker database.

### v3

@@ IV Conclusion
In this study, we developed and analyzed two hypergraph-based clustering algorithms, DIPHW and CoMem-DIPHW for improved cell clustering performance on scRNA-seq data. Firstly, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW) we developed extends the edge-dependent vertex weight (EDVW) hypergraph random walk proposed by [41], and consistently outperformed it in our experiments. DIPHW’s key idea lies in its symmetrized account of importance: it considers both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering. Secondly, the CoMem and CoMem-DIPHW clustering methods we developed incorporate a memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks to integrate both local and global information when computing embeddings for clustering. Thirdly, we identified and showed with experiments 3 the issue of inflated signals in co-expression networks brought by sparsity of scRNA-seq data, which provides additional motivation for the use of hypergraph representations. Fourthly, we compared our proposed methods with well-established clustering methods commonly used in scRNA-seq analysis packages. Our clustering performance evaluation spanned a broad range of modularity conditions, where CoMem-DIPHW and DIPHW consistently exhibited superior performance, particularly in weak modularity regimes. When clusters were harder to detect, baseline methods experienced a much more pronounced decay in performance compared to CoMem-DIPHW. For empirical analysis, we tested our algorithms on 4 real scRNA-seq datasets and 2 benchmark datasets. In these experiments, CoMem-DIPHW consistently demonstrated superior performance in differentiating and classifying distinct cell types compared to the baseline methods. Additionally, we designed a versatile simulation algorithm for scRNA-seq data that is highly customizable through its many user-specifiable parameters. This algorithm is designed to model a wide range of expression scenarios and modular structures. Notably, it incorporates intermodular co-expressions, which mimic cell-type cross-talk observed in real biological systems. For future work, we plan to optimize the implementation of CoMem-DIPHW to reduce its memory complexity, enhancing its scalability for larger datasets. We are also interested in exploring alternative ways to incorporate the memory mechanism into the hypergraph random walk. Currently, CoMem employs a simple product approach, moving beyond the simple product used in CoMem could potentially uncover even more effective clustering strategies.
@@
.1 Complexity Analysis
DIPHW
$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
CoMem-DIPHW
.2 Bipartite representation
.3 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
In Table S2, we evaluated how co-expression networks generated by different correlation computation methods lead to varying performance across various network-based clustering methods. The correlation methods considered were Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity. Our results show that cosine similarity often led to the worst clustering results among the correlation computation methods tested.
Method ARI NMI
Pearson Cosine Spearman Kendall Pearson Cosine Spearman Kendall
GreedyModularity 0.205 0.135 0.194 0.197 0.534 0.429 0.496 0.502
Louvain 0.493 0.322 0.445 0.473 0.784 0.666 0.768 0.781
Infomap 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Leiden -0.001 -0.001 -$0$0.000 0.062 0.040 0.082 0.073
Multilevel 0.032 0.034 0.034 0.035 0.305 0.289 0.325 0.319
Eigenvector 0.002 0.000 0.003 0.005 0.097 0.060 0.082 0.091
Table S2: Impact of Correlation Methods on Network-Based Clustering Performance. The table evaluates the impact of correlation computation methods—Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity—on the performance of various network-based clustering methods. The correlation computation method resulting in the worst clustering performance is marked in bold. Cosine similarity, which ignores shared zeros in cell expression profiles, often resulted in the worst clustering performance. These results show that, although using correlation computation methods that treat zeros the same as non-zero signals could inflate correlations, ignoring these zeros, as in cosine similarity, can lead to worse performance.
.4 Embedding Visualizations
@@
Figure S7: Clustering Performance of CoMem-DIPHW on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S8: Clustering Performance of PCA on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
@@
Figure S9: Clustering Performance of CoMem-DIPHW on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S10: Clustering Performance of PCA on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S11: Clustering Performance of CoMem-DIPHW on the Mouse Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database.
Figure S12: Clustering Performance of PCA on the Mouse Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. (c) Cell type annotation using the PanglaoDB marker database.